OpenStudy (aamey):

Find the equation of a sine wave that is obtained by shifting the graph of y=\sin(x) to the right 2 units and downward 9 units and is vertically stretched by a factor of 6 when compared to y=\sin(x).

OpenStudy (perl):

y = A sin( 2pi / period ( x - p.s.) ) + v.s.

OpenStudy (jdoe0001):

\(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Up\\ {\color{green}{ D}} < 0 & Down\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ A}} ( {\color{blue}{ B}}x + {\color{red}{ C}} ) + {\color{green}{ D}}\\ \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to \ left\\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to \ right\end{array} \end{array} }\) any ideas?

OpenStudy (jdoe0001):

the transformations are the same for say Asin(Bx+C) + D thus

OpenStudy (aamey):

@jdoe0001 is the period 2?

OpenStudy (jdoe0001):

period for sin(x) and cos(x) and is \(\large 2\pi\) if that's what you mean

OpenStudy (perl):

none of the transformations affected the period of the sine curve

OpenStudy (aamey):

@jdoe0001 I'm confused what to plug in

OpenStudy (jdoe0001):

yeah. you're not asked for a period just a horizontal shfit of 2 a vertical shift of 9 and a stretch of 6

OpenStudy (jdoe0001):

what you're asked is what would the value for "C" be and the value of "A" and the value of "D"

OpenStudy (aamey):

@jdoe0001 so is B 6?

OpenStudy (jdoe0001):

well. B would be 1 so \(\large \bf y = {\color{purple}{ A?}} sin( {\color{blue}{ 1}}x + {\color{red}{ C?}} ) + {\color{green}{ D?}}\)

OpenStudy (jdoe0001):

notice, is going 2 units to the right, so C is a negative, or C < 0 is also going downwards 9 units, so D is negative as well, or D < 0 and has a stretch factor of 6, so A should be obvious

OpenStudy (aamey):

@jdoe0001 since 2 is going to the right wouldn't that mean positive?

OpenStudy (jdoe0001):

one would expect that, yes but the graph behaviour is that, if it's negative, it goes to the right

OpenStudy (aamey):

@jdoe0001 I GOT IT THANK YOU SO SO MUCH!

OpenStudy (jdoe0001):

yw